# Efficient Reasoning for LLMs through Speculative Chain-of-Thought

### v1

@@ Abstract
Large reasoning language models such as OpenAI-o1 and Deepseek-R1 have recently attracted widespread attention due to their impressive task-solving abilities. However, the enormous model size and the generation of lengthy thought chains introduce significant reasoning costs and response latency. Existing methods for efficient reasoning mainly focus on reducing the number of model parameters or shortening the chain-of-thought length. In this paper, we introduce Speculative Chain-of-Thought (SCoT), which reduces reasoning latency from another perspective by accelerated average reasoning speed through large and small model collaboration. SCoT conducts thought-level drafting using a lightweight draft model. Then it selects the best CoT draft and corrects the error cases with the target model. The proposed thinking behavior alignment improves the efficiency of drafting and the draft selection strategy maintains the prediction accuracy for complex problems. Experimental results on GSM8K, MATH, GaoKao, CollegeMath and Olympiad datasets show that SCoT reduces reasoning latency by 48$\sim$66for Deepseek-R1-Distill-Qwen-32B while achieving near-target-model-level performance. Our code is available at https://github.com/Jikai0Wang/SpeculativeCoT.
@@ 1 Introduction
Large-scale reasoning language models (e.g., OpenAI-o1 (Jaech et al.,, 2024) and Deepseek-R1 (DeepSeek-AI et al.,, 2025)) have exhibited exceptional performance across a broad range of reasoning scenarios. Their superior performance stems from two key factors: massive model scale (e.g., DeepSeek-R1 with 671B parameters) and long chain-of-thought (CoT) (Wei et al.,, 2022) generation (reaching over 10,000 tokens for complex tasks). However, both factors incur substantial computational overhead, resulting in huge inference costs. In many reasoning systems where the model’s reasoning process remains implicit, excessive reasoning time induces significant response latency.
@@
Figure 1: Average reasoning accuracy and efficiency of Speculative Chain-of-Thought (SCoT) with Deepseek-R1-Distill-Qwen-32B and Deepseek-R1-Distill-Qwen-1.5B on GSM8K, MATH, GaoKao, CollegeMath and Olympiad. Graph (a): SCoT achieves near-target-model performance level on the 5 datasets. Graph (b): SCoT demonstrates an average speed-up ratio of 2.3× for reasoning latency. Graph (c): Comparison of reasoning efficiency-accuracy trade-off.
@@
In Section 3, we introduce thinking behavior alignment to improve drafting efficiency. We also design an error correction mechanism to preserve accuracy in dealing with complex tasks. We construct a training set by 500 samples from GSM8K for fine-tuning the target model to improve the capability of draft selection and error detection. We conduct comprehensive experiments in Section 4 to evaluate the effectiveness of SCoT. Figure 1 displays the overview reasoning accuracy and efficiency of SCoT. Adapting to task difficulty and draft quality, the draft selection mechanism decides when the target model should rethink. Simple tasks mainly utilize the draft model for thought chain generation, while complex problems engage the target model more extensively in the reasoning process. SCoT reduces reasoning latency by 48$\sim$66on 5 datasets (including the Olympiad-level dataset (He et al.,, 2024)) while maintaining near-target-model-level performance.
@@ 3.2 Generating Drafts of Thoughts
We use nuclear sampling for the small model thinking to enhance the diversity of drafts of CoTs. Thanks to their smaller parameter size, smaller models can draft reasoning chains significantly faster than large models when the length of the CoT is the same, thus speeding up the thinking process.
@@ 4.1 Settings
We adopt Deepseek-R1-Distill-Qwen-32B (DeepSeek-AI et al.,, 2025) (abbreviated as D-Qwen-32B) as the target model and Deepseek-R1-Distill-Qwen-1.5B as the draft model. We use a single A100-PCIE-40GB GPU for the 1.5B model and 4 A100-PCIE-40GB GPUs for the 32B model. We evaluate the efficiency and accuracy of Speculative CoT on reasoning datasets of various difficulty levels, including GSM8K (Cobbe et al.,, 2021), MATH (Hendrycks et al.,, 2021), GaoKao-En-2023 (Liao et al.,, 2024), CollegeMath (Tang et al.,, 2024) and Olympiad (He et al.,, 2024). As stated in Section 3, we use data from GSM8K training set to fine-tune the models with LoRA modules. Note that there is no overlap in the data used for training and evaluation. We set the rank of LoRA modules to 8 for both the draft and target model. The number of CoT drafts $n$ is set to 5. To avoid running out of memory when selecting the drafts, the maximum draft length is set to 5000. We compare the reasoning latency and accuracy of Speculative CoT with the vanilla reasoning by the 32B model.
@@ 4.2 Main Results
Table 1: Comparison of CoT length and reasoning efficiency. We report average CoT length of the target model $l_{\_}{M}$, average CoT length of the draft model $l_{\_}{M_{\_}{d}}$ and average CoT latency of each sample $t(s)$. Speed-up ratio $r$ is calculated by average thinking time compared with vanilla decoding.
@@
SCoT(32B+1.5B) 11.8 2.26 77.0 2.92 160.1 2.31 43.7 2.18 575.7 1.90
Lower Reasoning Latency  Table 1 demonstrates the average CoT length and the average CoT latency of each sample on the 5 datasets. $l_{\_}{M}$ and $l_{\_}{M_{\_}{d}}$ represent the numbers of tokens generated by the target model and the draft model during the thinking stage, respectively. Since the draft model generates 5 draft chains in parallel, $l_{\_}{M_{\_}{d}}$ refers to the length of the longest draft CoT. The average latency $t$ is measured in seconds. For the 32B model, $l_{\_}{M}$ exhibits a wide range of 302 to 8431 across different datasets. Except for MATH where the average CoT length of SCoT ($l_{\_}{M}+l_{\_}{M_{\_}{d}}$) is shorter than that of the 32B model, the overall CoT lengths of the two methods are close. For SCoT, most of the reasoning tokens are generated by the draft model, and the CoT tokens generated by the target model account for 8.8$\sim$45.9of the total CoT tokens. As task difficulty increases, small models exhibit higher error rates, thus necessitating greater involvement from the target model in the reasoning process. The draft selection mechanism dynamically determines whether the target model should reconsider its output, evaluating both problem complexity and draft generation quality. For simpler tasks, thought chains are predominantly produced by the draft model. Conversely, the system allocates more thought chain generation to the target model when handling complex problems.
The throughput of the 1.5B model is about 5 times that of the 32B model in our experimental environment. The CoT latency of SCoT is mainly composed of the target model reasoning time, draft model reasoning time, and draft selection time. For the GSM8K dataset, the proportions are approximately 33.9%, 60.2%, and 5.9% respectively. Generally, SCoT achieves greater speedup ratios on simpler tasks where the CoT drafts demonstrate higher quality. SCoT achieves a speedup ratio ranging from 1.90× to 2.92× compared to vanilla reasoning.
Near-target-model Level Accuracy  Figure 4 displays the accuracy of SCoT on these datasets. Comparative results indicate that SCoT maintains near-target-model performance levels on all datasets (even for the Olympiad-level task), with clearly superior accuracy compared to standalone draft model reasoning. For CollegeMath, SCoT even achieves a lossless acceleration.
Figure 4: Comparison of accuracy of the final answer on the 5 datasets.
@@ 4.3 Ablation Studies
Table 2: Comparative analysis of average CoT draft length on DeepSeek-R1-Distill-Qwen-1.5B with versus without thinking behavior alignment training.
@@
We conduct an ablation study in this section. We carry out 3 sets of experiments to observe the changes in accuracy on the 5 datasets with Deepseek-R1-Distill-Qwen: 1) SCoT with Single Draft: Only one CoT draft is generated for each sample (except for the special draft $T_{\_}{n+1}$), 2) SCoT without Tuning $M$: We use the original target model instead of the fine-tuned version to select the drafts, 3) SCoT without Error Correction: We remove $T_{\_}{n+1}$ and the rethinking strategy. Table LABEL:tab:ma shows the results. SCoT exhibits the highest accuracy on all datasets, indicating that both the three proposed techniques contribute to the reasoning performance.
In order to verify the effectiveness of thinking behavior alignment, we compare the average CoT draft length generated by the original Deepseek-R1-Distill-Qwen-1.5B and the fine-tuned version. Table 2 shows the mean CoT length on the 5 datasets. The CoT length generated by the fine-tuned model is significantly shorter than that of the original model on all datasets, especially for GSM8K. Therefore, thinking behavior alignment improves the efficiency of the drafting process.
@@ 4.5 Impact on Length of the Final Answer
We study the impact of SCoT on the length of the final answer in this section. Table 3 displays the answer length of the original 32B model and SCoT on the 5 datasets. In addition to reducing the latency of the reasoning phase, SCoT also reduces the length of the generated final answer compared to the target model and even brings additional inference efficiency improvements.

### v2

@@ Abstract
Large reasoning language models such as OpenAI-o1 and Deepseek-R1 have recently attracted widespread attention due to their impressive task-solving abilities. However, the enormous model size and the generation of lengthy thought chains introduce significant reasoning costs and response latency. Existing methods for efficient reasoning mainly focus on reducing the number of model parameters or shortening the chain-of-thought length. In this paper, we introduce Speculative Chain-of-Thought (SCoT), which reduces reasoning latency from another perspective by accelerated average reasoning speed through large and small model collaboration. SCoT conducts thought-level drafting using a lightweight draft model. Then it selects the best CoT draft and corrects the error cases with the target model. The proposed thinking behavior alignment improves the efficiency of drafting and the draft selection strategy maintains the prediction accuracy of the target model for complex tasks. Experimental results on GSM8K, MATH, GaoKao, CollegeMath and Olympiad datasets show that SCoT reduces reasoning latency by 48$\sim$66and 21$\sim$49for Deepseek-R1-Distill-Qwen-32B and Deepseek-R1-Distill-Llama-70B while achieving near-target-model-level performance. Our code is available at https://github.com/Jikai0Wang/SpeculativeCoT.
@@ 1 Introduction
Large-scale reasoning language models (e.g., OpenAI-o1 (Jaech et al.,, 2024) and Deepseek-R1 (DeepSeek-AI et al.,, 2025)) have exhibited exceptional performance across a broad range of reasoning scenarios. Their superior performance stems from two key factors: massive model scale (e.g., DeepSeek-R1 with 671B parameters) and long chain-of-thought (CoT) (Wei et al.,, 2023) generation (reaching over 10,000 tokens for complex tasks). However, both factors incur substantial computational overhead, resulting in huge inference costs. In many reasoning systems where the model’s reasoning process remains implicit, excessive reasoning time induces significant response latency.
@@
Figure 1: Average reasoning accuracy and efficiency of Speculative Chain-of-Thought (SCoT) with Deepseek-R1-Distill-Qwen-32B and Deepseek-R1-Distill-Llama-B on GSM8K, MATH, GaoKao, CollegeMath and Olympiad. Graph (a): SCoT achieves near-target-model performance level and outperforms TokenSkip on average generation accuracy on the 5 datasets. Graph (b): SCoT demonstrates an average speed-up ratio of 2.3× and 1.6× for the two target models on reasoning latency. Graph (c): Comparison of reasoning speed and accuracy for models of different sizes. Note that all the experiments for testing the reasoning speed are conducted on H800-80G GPUs. SCoT exhibits a better reasoning efficiency-accuracy trade-off through large and small model collaboration.
@@
In Section 3, we introduce thinking behavior alignment to improve drafting efficiency. We also design an error correction mechanism to preserve accuracy in dealing with complex tasks. We construct a training set by 500 samples from GSM8K to efficiently fine-tuning the target model to improve the capability of draft selection and error detection. We conduct comprehensive experiments in Section 4 to evaluate the effectiveness of SCoT. Figure 1 displays the overview reasoning accuracy and efficiency of SCoT. SCoT reduces reasoning latency by 48$\sim$66and 21$\sim$49for Deepseek-R1-Distill-Qwen-32B and Deepseek-R1-Distill-Llama-70B on 5 datasets (including the Olympiad-level dataset (He et al.,, 2024)) while maintaining near-target-model-level performance. It achieves better reasoning accuracy compared with TokenSkip (Xia et al.,, 2025) under similar speed-up ratios. Adapting to task difficulty and draft quality, the draft selection mechanism decides when the target model should rethink. Simple tasks mainly utilize the draft model for thought chain generation, while complex problems engage the target model more extensively in the reasoning process. On the other hand, the single-pass drafting and selection paradigm of SCoT significantly reduces the model reloading overhead between high-bandwidth memory and on-chip cache inherent in traditional speculative decoding methods, maintaining high efficiency even in long-sequence generation scenarios.
@@ 3.2 Generating Drafts of Thoughts
We use temperature sampling for the small model thinking to enhance the diversity of drafts of CoTs. Thanks to their smaller parameter size, smaller models can draft reasoning chains significantly faster than large models when the length of the CoT is the same, thus speeding up the thinking process.
@@ 4.1 Settings
We adopt Deepseek-R1-Distill-Qwen-32B (DeepSeek-AI et al.,, 2025) and Deepseek-R1-Distill-Llama-70B (abbreviated as D-Qwen-32B and D-Llama-70B) as the target models. We employ Deepseek-R1-Distill-Qwen-1.5B and Deepseek-R1-Distill-Llama-8B as the draft models for them, respectively. We conducte the two groups of experiments under different hardware environments. We use a single A100-PCIE-40GB GPU for the 1.5B model and 4 A100-PCIE-40GB GPUs for the 32B model. While we deploy the 8B model on a single H800-80GB GPU and the 80B model on 4 H800-80GB GPUs. We evaluate the efficiency and accuracy of Speculative CoT on reasoning datasets of various difficulty levels, including GSM8K (Cobbe et al.,, 2021), MATH (Hendrycks et al.,, 2021), GaoKao-En-2023 (Liao et al.,, 2024), CollegeMath (Tang et al.,, 2024) and Olympiad (He et al.,, 2024). As stated in Section 3, we use data from GSM8K training set to fine-tune the models with LoRA modules. Note that there is no overlap in the data used for training and evaluation.
For SCoT, we set the rank of LoRA modules to 8 for D-Qwen-1.5B/32B and D-Llama-8B, and we set it to 16 for D-Llama-70B. Learning rate is set to 5e-5 and 1e-4 for the draft models and target models, respectively. The number of CoT drafts $n$ is set to 5. We set the temperature for sampling to 0.6 in CoT drafting process. To avoid running out of memory when selecting the drafts, the maximum draft length is set to 5000. We set the maximum generation length of CoT to 20480. We compare the reasoning latency and accuracy of Speculative CoT with the vanilla reasoning by the target models.
We also compare the performance of SCoT with TokenSkip (Xia et al.,, 2025), which trains the target model with compressed CoT data to shorten the mean CoT length. The original CoT data are generated by the original target model. Then it compresses the thought chains through LLMLingua-2 (Pan et al.,, 2024). We adopt the training set of GSM8K to construct the training data. To compare the accuracy of the two methods at similar acceleration ratios, we set the compression ratio to 0.5 for TokenSkip, which means that about half of the tokens in thought chains are dropped.
@@ 4.2 Main Results
Table 1: Comparison of CoT length and reasoning efficiency. We report average CoT length of the target model $l_{\_}{M}$, average CoT length of the draft model $l_{\_}{M_{\_}{d}}$ and average CoT latency of each sample $t(\text{sec})$. Speed-up ratio $r$ is calculated by average thinking time compared with vanilla decoding.
@@
SCoT(32B+1.5B) 11.7 2.26 77.0 2.92 160.1 2.31 43.7 2.18 575.7 1.90
D-Llama-70B 21.6 1.00 105.5 1.00 188.3 1.00 63.3 1.00 443.6 1.00
SCoT(70B+8B) 11.0 1.96 57.8 1.83 111.0 1.70 50.4 1.26 325.0 1.36
Figure 4: Comparison of accuracy of the final answer on the 5 datasets with Deepseek-R1-Distill-Qwen (top) and Deepseek-R1-Distill-Llama (bottom). The compression ratio for TokenSkip is set to 0.5. The speed-up ratio in the legends represents the latency speed-up for TokenSkip and SCoT.
Lower Reasoning Latency  Table 1 demonstrates the average CoT length and the average CoT latency of each sample on the 5 datasets. $l_{\_}{M}$ and $l_{\_}{M_{\_}{d}}$ represent the numbers of tokens generated by the target model and the draft model during the thinking stage, respectively. Since the draft model generates 5 draft chains in parallel, $l_{\_}{M_{\_}{d}}$ refers to the length of the longest draft CoT. The average latency $t$ is measured in seconds. For the 32B model, $l_{\_}{M}$ exhibits a wide range of 302 to 8431 across different datasets. Except for MATH where the average CoT length of SCoT ($l_{\_}{M}+l_{\_}{M_{\_}{d}}$) is shorter than that of the 32B model, the overall CoT lengths of the two methods are close in the Qwen group. However, SCoT generates more tokens during the reasoning phase in the Llama group, where the speed-up ratio is relatively lower. For SCoT, most of the reasoning tokens are generated by the draft model, and the CoT tokens generated by the target model account for 2.8$\sim$45.9of the total CoT tokens. As task difficulty increases, small models exhibit higher error rates, thus necessitating greater involvement from the target model in the reasoning process. The draft selection mechanism dynamically determines whether the target model should reconsider its output, evaluating both problem complexity and draft generation quality. For simpler tasks, thought chains are predominantly produced by the draft model. Conversely, the proposed mechanism allocates more thought chain generation to the target model when handling complex problems.
The throughput of the draft model is about 5 times that of the corresponding target model for both groups in our experimental environment. The CoT latency of SCoT is mainly composed of the target model reasoning time, draft model reasoning time, and draft selection time. For the GSM8K dataset in the Qwen group, the proportions are approximately 33.9%, 60.2%, and 5.9% respectively. With longer generated CoTs (e.g., on Olympiad), draft selection incurs minimal overhead. Generally, SCoT achieves greater speed-up ratios on simpler tasks where the CoT drafts demonstrate higher quality. SCoT achieves a speed-up ratio ranging from 1.26× to 2.92× compared to vanilla reasoning.
Near-target-model Level Accuracy  Figure 4 displays the accuracy of different methods for Deepseek-R1-Distill-Qwen and Deepseek-R1-Distill-Llama on these datasets. Comparative results indicate that SCoT maintains near-target-model performance levels on all datasets (even for the Olympiad-level task), with clearly superior accuracy compared to standalone draft model reasoning. For CollegeMath, SCoT even achieves a better accuracy compared to the target model. Moreover, SCoT outperforms TokenSkip in generation accuracy on most groups except Olympiad with the Qwen model, even with higher speed-up ratios, exhibiting better accuracy-efficiency trade-off.
@@ 4.3 Ablation Studies
We conduct an ablation study in this section. We carry out 3 sets of experiments to observe the changes in accuracy on the 5 datasets with Deepseek-R1-Distill-Qwen: 1) SCoT with Single Draft: Only one CoT draft is generated for each sample (except for the special draft $T_{\_}{n+1}$), 2) SCoT without Tuning $M$: We use the original target model instead of the fine-tuned version to select the drafts, 3) SCoT without Error Correction: We remove $T_{\_}{n+1}$ and the rethinking strategy. Table LABEL:tab:ma shows the results. SCoT exhibits the highest accuracy on all datasets, indicating that both of the three proposed techniques contribute to the reasoning performance. Drafting multiple thought chains increases the upper limit of draft quality, thus enhancing the potential for higher-quality outputs. In preliminary experiments, we found that without fine-tuning, $M$ never selects the “All reasoning paths are wrong.” option for any CoT draft candidates. After training, the target model can acquire a certain error correction capability. We will discuss this further in Section 4.5.
Table 2: Comparative analysis of average CoT draft length on DeepSeek-R1-Distill-Qwen-1.5B and Deepseek-R1-Distill-Llama-8B with versus without thinking behavior alignment training.
@@
Deepseek-R1-Distill-Llama-8B 585 1979 3223 1285 6474
+ Thinking Behavior Alignment 258 1536 2253 1254 3783
In order to verify the effectiveness of thinking behavior alignment, we compare the average CoT draft length generated by the original draft models and the fine-tuned versions. Table 2 shows the mean CoT length on the 5 datasets. The CoT length generated by the fine-tuned model is significantly shorter than that of the original model on all datasets for both the two draft models, especially for GSM8K. Therefore, thinking behavior alignment improves the efficiency of the drafting process.
4.4 Comparison with Speculative Decoding
Table 3: Comparison of reasoning efficiency with auto-regressive decoding (ARD) and speculative decoding (SD) on the five datasets with the two models. $s$ represents the average throughput during the reasoning phase (including the prefilling phase). For SCoT, only valid tokens (generated by either the draft model or the target model) are included in the calculation of $s$. $r^{\prime}$ represents the speed-up ratio of throughput. $r$ stands for the speed-up ratio of average reasoning latency of each sample.
Method GSM8K MATH GaoKao CollegeMath Olympiad
$s$ $r^{\prime}$ $r$ $s$ $r^{\prime}$ $r$ $s$ $r^{\prime}$ $r$ $s$ $r^{\prime}$ $r$ $s$ $r^{\prime}$ $r$
Deepseek-R1-Distill- Qwen-32B (A100-PCIE-40GB)
ARD 11.4 1.00 1.00 9.4 1.00 1.00 8.7 1.00 1.00 11.0 1.00 1.00 7.9 1.00 1.00
SD 18.5 1.63 1.87 13.4 1.42 1.41 12.1 1.38 1.32 15.6 1.42 1.34 9.1 1.16 0.59
SCoT 23.9 2.10 2.26 25.4 2.69 2.92 19.1 2.18 2.31 27.9 2.53 2.18 11.2 1.42 1.90
Deepseek-R1-Distill- Llama-70B (H800-80GB)
ARD 14.5 1.00 1.00 14.1 1.00 1.00 13.4 1.00 1.00 14.5 1.00 1.00 12.7 1.00 1.00
SD 25.1 1.64 1.63 21.2 1.50 1.49 17.6 1.31 1.35 23.2 1.59 1.61 12.6 0.99 0.58
SCoT 37.3 2.45 1.96 41.8 2.96 1.83 28.7 2.14 1.70 37.5 2.58 1.26 17.4 1.37 1.36
We compare the reasoning efficiency of SCoT with speculative decoding with large and small models collaboration (Leviathan et al.,, 2023; Kim et al.,, 2023) in this section. Speculative decoding generates several tokens with the draft model and then verifies the candidates with the target model in each decoding step. Tokens that match the original output will be accepted during verification. We evaluate the throughput and latency with Deepseek-R1-Distill- Qwen-32B and Deepseek-R1-Distill-Llama-70B. We use the original Deepseek-R1-Distill-Qwen-1.5B and Deepseek-R1-Distill-Llama-8B as the draft models for speculative decoding. The maximum draft length is set to 5.
Table 3 displays the results. The throughput of auto-regressive decoding is affected by the length of the sequence. Speculative decoding performs well when the sequence length is short. It achieves a reasoning latency speed-up ratio of up to 1.87 and 1.63 for the Qwen model and the Llama model, respectively. However, for the Olympiad dataset, where the average generation length exceeds 5000 tokens, the drafting overhead grows with sequence length, ultimately causing deceleration for both the target models. Note that speculative decoding may produce outputs that deviate from the original model when using half-precision. In certain cases (e.g., GSM8K dataset with the Qwen model), the shorter generation lengths can lead to $r^{\prime}<r$. Employing the same draft models, SCoT achieves a higher throughput on all datasets. It shows a speed-up ratio of up to 2.96 for throughput. It also exhibits a better acceleration for reasoning latency (except for CollegeMath in the Llama group). While speculative decoding requires multiple verifications, SCoT’s single draft-verify cycle significantly reduces the frequent model reloads between GPU high-bandwidth memory (HBM) and on-chip cache. Therefore, generation length exhibits a relatively minor impact on SCoT’s efficiency (as stated in Section 4.2 the speed-up ratio of SCoT is mainly affected by the difficulty of the task), reflecting its good adaptability to long sequence generation scenarios.
@@ 4.6 Impact on Length of the Final Answer
We study the impact of SCoT on the length of the final answer in this section. Table 4 displays the answer length of SCoT on the 5 datasets. In addition to reducing the latency of the reasoning phase, SCoT also reduces the length of the generated final answer compared to the target model and even brings additional inference efficiency improvements in the Qwen group. For the Llama group, SCoT generates final answers with lengths comparable to those produced by the original target model.
@@
Deepseek-R1-Distill-Llama-70B 238 256 453 312 840
SCoT (70B+8B) 212 310 568 314 1068
4.7 Performance under Different Number of Chains
Figure 5: Impact of the chain number on reasoning accuracy and latency on GSM8K.
To explore the best number of CoT draft candidates, we evaluate the reasoning accuracy and latency under different numbers of draft thought chains in this section. We conduct an experiment with the Deepseek-R1-Distill-Qwen models on the GSM8K dataset. We tested the performance when drafting 1 to 8 thought chains. Figure 5 displays the results. The blue and green lines represent the accuracy and reasoning latency, respectively. The two dotted lines show the baseline accuracy and latency of Deepseek-R1-Distill-Qwen-32B. Increasing the number of chains increases the probability of including high-quality reasoning paths in the set of drafts. When using 2 draft chains, while the latency reduction remains modest, SCoT surprisingly outperforms the target model in accuracy, demonstrating the remarkable potential of collaborative model generation. A higher speed-up ratio indicates that SCoT accepts more reasoning chains generated by the draft model. SCoT can reduce its reliance on the target model due to the increase in the number of candidates, thus achieving a higher speed-up ratio. However, as the number of chains increases, the difficulty of training the target model to select the correct draft also increases. Meanwhile, due to the inherent randomness in training procedures and the limited scale of training data, it exhibits fluctuations in reasoning accuracy. Moreover, an oversupply of chains complicates error detection for the target model, which may lead to accuracy degradation despite the improved speed. Overall, SCoT demonstrates optimal acceleration performance with 5 thought chains while maintaining acceptable accuracy degradation.